# ELIP: Enhanced Visual-Language Foundation Models for Image Retrieval

### v1

@@ 7.2 Results on OOD Benchmarks
The results on the out-of-distribution benchmarks are given in Table 3. It can be observed that our method achieves a zero-shot improvement when tested on the OOD benchmarks Occluded COCO and ImageNet-R, demonstrating the generalisation capability of the ELIP models. After fine-tuning on COCO and ImageNet (not the OOD version), the performance of the models is further boosted, demonstrating the effectiveness of our framework. Please refer to appendix for details on the fine-tuning.
@@ 8 Conclusion
In this paper, we propose Enhance Language-Image Pre-training (ELIP) to enhance visual-language foundation models for text-based image retrieval, and have achieved significant improvement on CLIP/SigLIP as well as outperformed the state-of-the-art BLIP-2 model.
Our architecture, a simple MLP mapping network, generates a set of visual prompt vectors which improves the selection of image features relevant to the text query, therefore boosting the performance of visual-language foundation models.
We have also explored a ‘student friendly’ best practice to train our architecture with limited resources, which involves grouping hard samples in training batches and selection of training data with maximum information.
Furthermore, in complementary to existing text-to-image retrieval benchmarks, we have set up two new benchmarks to evaluate different models’ ability in out-of-distribution scenarios and observe our ELIP models have outperformed CLIP and BLIP-2.
@@ A.1 Choice of $k$ for re-ranking
Table 4: Recall @ Different $k$ for CLIP and BLIP-2 initial ranking. See text for a more detailed discussion.
In terms of the selection of the top-$k$ values for re-ranking as mentioned in Section 7, the value of $k$ is chosen such that the recall at that $k$ is high in the original ranking, and also that the inference is fast. More specifically, as in Table 4, for ELIP-C, the recall@100 for COCO/Flickr is over 95%, recall@500 for Occluded COCO is over 80%, recall@1000 for ImageNet-R is over 95%; for ELIP-B, the recall@20 for COCO/Flickr is over 95%, recall@100 for Occluded COCO is over 75%, recall@200 for ImageNet-R is over 80%. We have experimented with increasing $k$ for re-ranking, but there is no significant performance boost while the computational cost is much higher, so we keep our choice of the $k$ values as a trade off between performance and computational efficiency.
@@ A.2 Implementation Details for General Experiments
As mentioned in Section 5.1, in practice we use a random subset of 6M samples for training ELIP-C/ELIP-S and a random subset of 1M samples for training ELIP-B. For ELIP-B, we first select the samples with the top 10% highest learnability as described in Section 5.2, and then train the ELIP-B architecture with the selected training samples.
@@ A.3 Implementation Details for Fine-tuning on COCO and ImageNet
As mentioned in Section 7.2, we have further fine-tuned our models (pre-trained on DataCompDR) on COCO and ImageNet. For both datasets, we first group the hard samples in batches using the method described in Section 5.1. The captions for the images are category names, and we ensure that in a batch one category can only appear once, so that there will not be noisy in training. We fine-tune ELIP-C and ELIP-B with an initial learning rate of $51e-$ 5 for 10k iterations. Although there is still a domain gap between COCO and Occluded COCO, ImageNet and ImageNet-R, we find that even fine-tuning on COCO and ImageNet can already bring a boost to the performance on Occluded COCO and ImageNet-R as the caption domain is closer.
@@ Appendix B Additional Qualitative Results
In this section, we provide more qualitative results as mentioned in Section 7 of the main paper, including: 1) Qualitative comparison between the rankings of CLIP and ELIP-C on COCO (Figure 9), Flickr (Figure 10), Occluded COCO (Figure 11) and ImageNet-R (Figure 12); 2) Qualitative comparison between the rankings of BLIP-2 and ELIP-B on COCO (Figure 13), Flickr (Figure 14), Occluded COCO (Figure 15) and ImageNet-R (Figure 16); 3) Attention maps for CLIP and ELIP-C (Figure 17 and Figure 18); 4) Attention maps for BLIP-2 and ELIP-B (Figure 19 and Figure 20).
@@ B.1 Qualitative Comparison of CLIP and ELIP-C
Figure 9 shows more qualitative comparison between the CLIP initial ranking and ELIP-C re-ranking on COCO. For each example, we show both top-5 rankings (from left to right) and highlight the ground truth image in black dashed box.
In the first and second examples (from top to bottom), the text queries are ‘Two young men playing a game of soccer’, and ‘A man cuts into a small cake with his sharp knife’, and the ground truth images are not in top-5 images of the CLIP initial ranking, but ranked top-1 in our ELIP-C re-ranking.
Figure 9: Qualitative Comparison on COCO for CLIP and ELIP-C. Please see text for a more detailed discussion.
Figure 10 shows more qualitative comparison between the CLIP initial ranking and ELIP-C re-ranking on Flickr. For each example, we show both top-5 rankings (from left to right) and highlight the ground truth image in black dashed box.
In the first example, the text query is ‘A man in blue overalls and red shirt holding a chainsaw’, and the ground truth image ranks top-3 in the initial ranking but top-1 in our re-ranking. In the second example, the text query is ‘A muzzled greyhound dog wearing yellow and black is running on the track’, and the ground truth image is not in top-5 images of the CLIP initial ranking, but ranked top-1 in our ELIP-C re-ranking.
Figure 10: Qualitative Comparison on Flickr for CLIP and ELIP-C. Please see text for a more detailed discussion.
Figure 11 displays qualitative comparison of CLIP initial ranking and our ELIP-C re-ranking on Occluded COCO. We display the top-5 retrieved images. Negative samples (errors) are highlighted in orange solid box.
For the first example, the text query is ‘bowl’ and CLIP confuses it with ‘frisbee’, ‘toilet’ and ‘plate’ in top-5 retrieved images; for the second example, the text query is ‘boat’ and CLIP confuses it with ‘surfboard’ in top-5 retrieved images.
Figure 11: Qualitative Comparison on Occluded COCO for CLIP and ELIP-C. Please see text for a more detailed discussion.
Figure 12 displays qualitative comparison of CLIP initial ranking and our ELIP-C re-ranking on ImageNet-R. We display examples at the top-100 rankings where there is difference between the two models, i.e., one model retrieves a positive sample while the other model retrieves a negative sample, as the majority of the top-100 retrieves samples are positive. Generally, ELIP-C retrieves more positive samples in top-100 images than CLIP. Negative samples (errors) are highlighted in orange solid box.
For the first example, the text query is ‘hen’ and CLIP retrieves some ‘duck’s in top-100 retrieved images; for the second example, the text query is ‘baseball player’ and CLIP retrieves some players for other ball games in top-100 retrieved images while our ELIP-C retrieves a basketball player in top-100 images.
Figure 12: Qualitative Comparison on ImageNet-R for CLIP and ELIP-C. Please see text for a more detailed discussion.
@@
B.2 Qualitative Comparison of BLIP and ELIP-B
@@
Figure 13 shows more qualitative comparison between the BLIP-2 ranking and ELIP-B ranking on COCO. For each example, we show both top-5 rankings (from left to right) and highlight the ground truth image in black dashed box.
In the first and second examples, the text queries are ‘There is dust coming out of the catcher’s glove as a boy prepares to bat’ and ‘Park scene, park bench, light pole and building in background, probably city park area’, and the ground truth images are ranked top-5 / top-4 respectively by BLIP-2, but ranked top-1 by ELIP-B.
Figure 13: Qualitative Comparison on COCO for BLIP-2 and ELIP-B. Please see text for a more detailed discussion.
@@
Figure 14 shows more qualitative comparison between the BLIP-2 ranking and ELIP-B ranking on Flickr. For each example, we show both top-5 rankings (from left to right) and highlight the ground truth image in black dashed box.
In the first and second examples, the text queries are ‘A brown dog waits for the frisbee to come down before catching it’ and ‘A crowd is present at a bar’, and the ground truth images rank top-2 in the BLIP-2 ranking but top-1 in our ELIP-B ranking.
Figure 14: Qualitative Comparison on Flickr for BLIP-2 and ELIP-B. Please see text for a more detailed discussion.
@@
Figure 15 displays qualitative comparison of BLIP-2 initial ranking and our ELIP-B re-ranking on Occluded COCO. Similar to Figure 12, we display examples at the top-100 rankings where there is a difference between the two models, i.e., one model retrieves a positive sample while the other model retrieves a negative sample. Generally, ELIP-B retrieves more positive samples in top-100 images than BLIP-2. Negative samples (errors) are highlighted in orange solid box.
For the first example, the text query is ‘cell phone’ and BLIP-2 confuses it with ‘remote control’ and ‘camera’ in top-100 retrieved images, while ELIP-B confuses it with ‘keyboard’; for the second example, the text query is ‘knife’ and BLIP-2 confuses it with ‘scissors’, ‘fruit peeler’ and ‘pizza cutter’ in top-100 retrieved images, while ELIP-B confuses it with ‘scissors’.
Figure 15: Qualitative Comparison on Occluded COCO for BLIP-2 and ELIP-B. Please see text for a more detailed discussion.
@@
Figure 16 displays qualitative comparison of BLIP-2 initial ranking and our ELIP-B re-ranking on ImageNet-R. Similarly, we display examples at the top-100 rankings where there is a difference between the two models, i.e., one model retrieves a positive sample while the other model retrieves a negative sample. Generally, ELIP-B retrieves more positive samples in top-100 images than BLIP-2. Negative samples (errors) are highlighted in orange solid box.
For the third example, the text query is ‘hammerhead’ and BLIP-2 retrieves some ‘hammer’s in top-100 retrieved images; for the fourth example, the text query is ‘mitten’ and BLIP-2 retrieves some other objects made of similar material of mitten in top-100 retrieved images while our ELIP-B makes a similar mistake but the total number of errors are lower.
Figure 16: Qualitative Comparison on ImageNet-R for BLIP-2 and ELIP-B. Please see text for a more detailed discussion.
@@
In conclusion, these results demonstrate that, our ELIP-B re-ranking model further improves the discrimination capability between the ground truth and hard negative images, achieving a better retrieval performance than the original BLIP-2 model.
B.3 Attention Map of CLIP and ELIP-C

### v2

@@ 7.2 Results on OOD Benchmarks
The results on the out-of-distribution (OOD) benchmarks are presented in Table 3. ELIP achieves notable zero-shot improvements across all models on the OOD benchmarks, Occluded COCO and ImageNet-R, highlighting the strong generalization capabilities of the ELIP models. The performance can be improved further by fine-tuning the mapping network on suitable datasets (the image and text encoders are frozen). Since it is not feasible to fine-tune on Occluded COCO (very few data samples) and ImageNet-R (evaluation only), for Occluded COCO retrieval we fine-tune on the original COCO dataset, and for ImageNet-R retrieval we fine-tune on ImageNet. As can be seen in Table 3 by this fine-tuning the performance of all the models is significantly boosted further. This demonstrates that fine-tuning ELIP enables efficient adaptation of the models to new datasets. The significant difference ELIP makes is also illustrated in Figure 6 (left). Please refer to the appendix on the fine-tuning.
@@ 8 Conclusion
In this paper, we introduced Enhance Language-Image Pre-training (ELIP), a method to improve visual-language foundation models for text-based image retrieval. ELIP is a simple plug-and-play modification to pre-trained visual-language foundation models that significantly improves their zero-shot performance. Furthermore, the mapping network can be fine-tuned to efficiently adapt these models to OOD datasets, leading to still further improvements. We have also demonstrated, by visualizing the attention maps, that ELIP enables the image encoder to attend to more relevant details.
Future work could apply ideas similar to ELIP to enhance generative Multimodal Large Language Models by introducing more effective text-guided visual attention and encoding for both decoder-only architectures [54] and cross-attention-based architectures [3].
@@ A.1 Choice of $k$ for re-ranking
SigLIP
Dataset Top-20 Top-50 Top-100 Top-200 Top-500
COCO 90.14 96.07 98.77 99.44 99.78
Flickr 98.98 99.64 99.82 99.92 100.00
Dataset Top-50 Top-100 Top-200 Top-500 Top-1000
Occluded COCO 68.84 75.90 81.18 86.96 91.94
ImageNet-R 39.22 67.46 90.46 98.14 99.04
SigLIP-2
Dataset Top-20 Top-50 Top-100 Top-200 Top-500
COCO 91.10 96.59 98.85 99.55 99.80
Flickr 99.34 99.80 99.92 99.94 100.00
Dataset Top-50 Top-100 Top-200 Top-500 Top-1000
Occluded COCO 70.69 79.57 85.67 91.28 95.34
ImageNet-R 39.39 67.74 90.90 98.24 99.10
@@
Table 4: Recall @ Different $k$ for CLIP, SigLIP and BLIP-2 initial ranking. See text for a more detailed discussion.
In terms of the selection of the top-$k$ values for re-ranking as mentioned in Section 7, the value of $k$ is chosen such that the recall at that $k$ is high in the original ranking, and also so that the inference is fast. For ELIP-C, $k$ is set to 100 for COCO and Flickr, 500 for Occluded COCO, and 1000 for ImageNet-R; for ELIP-S and ELIP-S-2, $k$ is set to 100 for COCO and Flickr, 500 for Occluded COCO, and 200 for ImageNet-R; for ELIP-B, $k$ is set to 20 for COCO and Flickr, 100 for Occluded COCO, and 200 for ImageNet-R. More specifically, as in Table 4, for CLIP, the recall@100 for COCO/Flickr is over 95%, recall@500 for Occluded COCO is over 80%, recall@1000 for ImageNet-R is over 95%; For SigLIP/SigLIP-2, the recall@100 for COCO/Flickr is over 95%, recall@500 for Occluded COCO is over 85%, recall@200 for ImageNet-R is over 90%; for BLIP-2, the recall@20 for COCO/Flickr is over 95%, recall@100 for Occluded COCO is over 75%, recall@200 for ImageNet-R is over 80%. We have experimented with increasing $k$ for re-ranking, but there is no significant performance boost while the computational cost is much higher.
@@ A.2 Implementation Details for General Experiments
As mentioned in Section 5.1, in practice we use a random subset of 6M samples for training ELIP-C/ELIP-S and a random subset of 1M samples for training ELIP-B. For ELIP-B, we first select the samples with the top 10% highest learnability as described in Section 5.2, and then train the ELIP-B architecture with the selected training samples. For ELIP-C, we use the default backbone (ViT-B) given limited computing resources and for the ablations, while for ELIP-S and ELIP-B we use the best backbone (ViT-G for BLIP-2 and SigLIP-2, ViT-SO400M for SigLIP) to show the effectiveness of our method over the state-of-the-art model. As we use large backbones for ELIP-S and it increases GPU memory, for every training batch we take 10 samples out of the batch of 40 in the ELIP-C training dataset.
The implementation of the MLP Mapping Network is three linear layers with a GELU between each pair of layers. We expand the output dimension to be $n$ times when we generate $n$ tokens and then divide the generated vector into $n$ tokens. We have also investigated other architectures for the Mapping Network, such as transformer layers, different numbers of layers of MLP, and linear network, but there is no significant performance advantage found. We therefore use the current architecture which is simple but effective.
@@ A.3 Implementation Details for Fine-Tuning on COCO and ImageNet
As mentioned in Section 7.2, we have further fine-tuned our models (pre-trained on DataCompDR) on COCO and ImageNet, respectively. For both datasets, we first group the hard samples in batches using the method described in Section 5.1. The captions for the images are category names, and we ensure that in a batch one category can only appear once, so that there will not be noise in training. We fine-tune ELIP-C, ELIP-S, and ELIP-B with an initial learning rate of $51e-$ 5 for 10k iterations. Although there is still a domain gap between COCO and Occluded COCO, ImageNet, and ImageNet-R, we find that even fine-tuning on COCO and ImageNet can already bring a boost to the performance on Occluded COCO and ImageNet-R as the caption domain is closer. It is not feasible to fine-tune on Occluded COCO and ImageNet-R because Occluded COCO has very few data samples and ImageNet-R is designed for evaluation.
@@ Appendix B Additional Qualitative Results
In this section, we provide more qualitative results as mentioned in Section 7 of the main paper, including: 1) Qualitative comparison between the rankings of CLIP and ELIP-C on COCO (Figure 9), Flickr (Figure 10), Occluded COCO (Figure 11) and ImageNet-R (Figure 12); 2) Qualitative comparison between the rankings of SigLIP/SigLIP-2 and ELIP-S/ELIP-S-2 on COCO (Figure 13), Flickr (Figure 14), Occluded COCO (Figure 15) and ImageNet-R (Figure 16); 3) Qualitative comparison between the rankings of BLIP-2 and ELIP-B on COCO (Figure 17), Flickr (Figure 18), Occluded COCO (Figure 19) and ImageNet-R (Figure 20); 4) Attention maps for CLIP and ELIP-C (Figure 21 and Figure 22); 5) Attention maps for SigLIP/SigLIP-2 and ELIP-S/ELIP-S-2(Figure 23 and Figure 24); 6) Attention maps for BLIP-2 and ELIP-B (Figure 25 and Figure 26).
@@ B.1 Qualitative Comparison of CLIP and ELIP-C
Figure 9 shows more qualitative comparison between the CLIP initial ranking and ELIP-C re-ranking on COCO.
Figure 9: Qualitative Comparison on COCO for CLIP and ELIP-C. For each example, we show both top-5 rankings (from left to right) and highlight the ground truth image in a black dashed box. In the first and second examples (from top to bottom), the text queries are ‘Two young men playing a game of soccer’, and ‘A man cuts into a small cake with his sharp knife’, and the ground truth images are not in top-5 images of the CLIP initial ranking, but ranked top-1 in our ELIP-C re-ranking.
Figure 10 shows more qualitative comparison between the CLIP initial ranking and ELIP-C re-ranking on Flickr.
Figure 10: Qualitative Comparison on Flickr for CLIP and ELIP-C. For each example, we show both top-5 rankings (from left to right) and highlight the ground truth image in a black dashed box. In the first example, the text query is ‘A man in blue overalls and red shirt holding a chainsaw’, and the ground truth image ranks top-3 in the initial ranking but top-1 in our re-ranking. In the second example, the text query is ‘A muzzled greyhound dog wearing yellow and black is running on the track’, and the ground truth image is not in top-5 images of the CLIP initial ranking, but ranked top-1 in our ELIP-C re-ranking.
Figure 11 displays qualitative comparison of CLIP initial ranking and our ELIP-C re-ranking on Occluded COCO.
Figure 11: Qualitative Comparison on Occluded COCO for CLIP and ELIP-C. We display the top-5 retrieved images. Negative samples (errors) are highlighted in an orange solid box. For the first example, the text query is ‘bowl’ and CLIP confuses it with ‘frisbee’, ‘toilet’, and ‘plate’ in top-5 retrieved images; for the second example, the text query is ‘boat’ and CLIP confuses it with ‘surfboard’ in top-5 retrieved images.
Figure 12 displays qualitative comparison of CLIP initial ranking and our ELIP-C re-ranking on ImageNet-R.
Figure 12: Qualitative Comparison on ImageNet-R for CLIP and ELIP-C. We display examples at the top-100 rankings where there is a difference between the two models, i.e., one model retrieves a positive sample while the other model retrieves a negative sample, as the majority of the top-100 retrieves samples are positive. Generally, ELIP-C retrieves more positive samples in top-100 images than CLIP. Negative samples (errors) are highlighted in an orange solid box. For the first example, the text query is ‘hen’ and CLIP retrieves some ‘duck’s in top-100 retrieved images; for the second example, the text query is ‘baseball player’ and CLIP retrieves some players for other ball games in the top-100 retrieved images while our ELIP-C retrieves a basketball player in top-100 images.
@@
B.2 Qualitative Comparison of SigLIP/SigLIP-2 and ELIP-S/ELIP-S-2
Figure 13 shows more qualitative comparison between the SigLIP/SigLIP-2 initial ranking and ELIP-S/ELIP-S-2 re-ranking on COCO.
@@ B.2 Qualitative Comparison of SigLIP/SigLIP-2 and ELIP-S/ELIP-S-2
Figure 13: Qualitative Comparison on COCO for SigLIP and ELIP-S (top), SigLIP-2 and ELIP-S-2 (bottom). For each example, we show both top-5 rankings (from left to right) and highlight the ground truth image in a black dashed box. In the first (SigLIP v.s. ELIP-S) and second (SigLIP-2 v.s. ELIP-S-2) examples (from top to bottom), the text queries are ‘A small blue plane sitting on top of a field’, and ‘Assortment of doughnuts and other snack items on a serving tray’, and the ground truth images are not in top-5 images of the SigLIP/SigLIP-2 initial ranking, but ranked top-1 in our ELIP-S/ELIP-S-2 re-ranking.
Figure 14 shows more qualitative comparison between the SigLIP/SigLIP-2 initial ranking and ELIP-S/ELIP-S-2 re-ranking on Flickr.
@@
Figure 14: Qualitative Comparison on Flickr for SigLIP and ELIP-S (top), SigLIP-2 and ELIP-S-2 (bottom). For each example, we show both top-5 rankings (from left to right) and highlight the ground truth image in a black dashed box. In the first example, the text query is ‘A man wearing bathing trunks is parasailing in the water’, and the ground truth image ranks top-3 in the SigLIP initial ranking but top-1 in our ELIP-S re-ranking. In the second example, the text query is ‘A group of mountain bikers race each other down a dirt hill’, and the ground truth image ranks top-3 in the SigLIP-2 initial ranking, but ranked top-1 in our ELIP-S-2 re-ranking.
Figure 15 displays qualitative comparison of SigLIP/SigLIP-2 initial ranking and our ELIP-S/ELIP-S-2 re-ranking on Occluded COCO.
@@
Figure 15: Qualitative Comparison on Occluded COCO for SigLIP and ELIP-S (top), SigLIP-2 and ELIP-S-2 (bottom). Similar to Figure 12, we display examples at the top-100 rankings where there is a difference between the two models, i.e., one model retrieves a positive sample while the other model retrieves a negative sample. Generally, ELIP-S/ELIP-S-2 retrieves more positive samples in top-100 images than SigLIP/SigLIP-2. Negative samples (errors) are highlighted in an orange solid box. For the first example, the text query is ‘bowl’ and SigLIP confuses it with ‘roller skating rink’, ‘plate’, and ‘frisbee’ in the top-100 retrieved images, while ELIP-S confuses it with ‘cup’; for the second example, the text query is ‘TV’ and SigLIP-2 confuses it with ‘cell phone’, ’laptop’ and ’remote control’ in top-100 retrieved images.
Figure 16 displays qualitative comparison of SigLIP/SigLIP-2 initial ranking and our ELIP-S/ELIP-S-2 re-ranking on ImageNet-R.
@@
Figure 16: Qualitative Comparison on ImageNet-R for SigLIP and ELIP-S (top), SigLIP-2 and ELIP-S-2 (bottom). We display examples at the top-100 rankings where there is a difference between the two models, i.e., one model retrieves a positive sample while the other model retrieves a negative sample, as the majority of the top-100 retrieves samples are positive. Generally, ELIP-S/ELIP-S-2 retrieves more positive samples in top-100 images than SigLIP/SigLIP-2. Negative samples (errors) are highlighted in an orange solid box. For the first example, the text query is ‘duck’ and SigLIP retrieves some ‘goose’s in top-100 retrieved images while our ELIP-S retrieves a ‘goose’ in top-100 images; for the second example, the text query is ‘centipede’ and SigLIP-2 retrieves some other insects in top-100 retrieved images.
@@
In conclusion, these results demonstrate that, though the initial SigLIP/SigLIP-2 model can roughly retrieve relevant images, it is less able to distinguish fine-grained differences between the top retrieved images, whereas our re-ranking model further improves the discrimination between the ground truth and hard negative images, achieving a better retrieval.
B.3 Qualitative Comparison of BLIP and ELIP-B
Figure 17 shows more qualitative comparison between the BLIP-2 ranking and ELIP-B re-ranking on COCO.
Figure 17: Qualitative Comparison on COCO for BLIP-2 and ELIP-B. For each example, we show both top-5 rankings (from left to right) and highlight the ground truth image in a black dashed box. In the first and second examples, the text queries are ‘There is dust coming out of the catcher’s glove as a boy prepares to bat’ and ‘Park scene, park bench, light pole and building in background, probably city park area’, and the ground truth images are ranked top-5 / top-4 respectively by BLIP-2, but ranked top-1 by ELIP-B.
Figure 18 shows more qualitative comparison between the BLIP-2 ranking and ELIP-B re-ranking on Flickr.
Figure 18: Qualitative Comparison on Flickr for BLIP-2 and ELIP-B. For each example, we show both top-5 rankings (from left to right) and highlight the ground truth image in a black dashed box. In the first and second examples, the text queries are ‘A brown dog waits for the frisbee to come down before catching it’ and ‘A crowd is present at a bar’, and the ground truth images rank top-2 in the BLIP-2 ranking but top-1 in our ELIP-B ranking.
Figure 19 displays qualitative comparison of BLIP-2 initial ranking and our ELIP-B re-ranking on Occluded COCO.
Figure 19: Qualitative Comparison on Occluded COCO for BLIP-2 and ELIP-B. Similar to Figure 12, we display examples at the top-100 rankings where there is a difference between the two models, i.e., one model retrieves a positive sample while the other model retrieves a negative sample. Generally, ELIP-B retrieves more positive samples in top-100 images than BLIP-2. Negative samples (errors) are highlighted in an orange solid box. For the first example, the text query is ‘cell phone’ and BLIP-2 confuses it with ‘remote control’ and ‘camera’ in the top-100 retrieved images, while ELIP-B confuses it with ‘keyboard’; for the second example, the text query is ‘knife’ and BLIP-2 confuses it with ‘scissors’, ‘fruit peeler’ and ‘pizza cutter’ in top-100 retrieved images, while ELIP-B confuses it with ‘scissors’.
Figure 20 displays qualitative comparison of BLIP-2 initial ranking and our ELIP-B re-ranking on ImageNet-R.
Figure 20: Qualitative Comparison on ImageNet-R for BLIP-2 and ELIP-B. Similarly, we display examples at the top-100 rankings where there is a difference between the two models, i.e., one model retrieves a positive sample while the other model retrieves a negative sample. Generally, ELIP-B retrieves more positive samples in top-100 images than BLIP-2. Negative samples (errors) are highlighted in an orange solid box. For the third example, the text query is ‘hammerhead’ and BLIP-2 retrieves some ‘hammer’s in the top-100 retrieved images; for the fourth example, the text query is ‘mitten’ and BLIP-2 retrieves some other objects made of similar material of mitten in top-100 retrieved images while our ELIP-B makes a similar mistake but the total number of errors are lower.
In conclusion, these results demonstrate that the ELIP-B re-ranking model improves the discrimination capability between the positive and hard negative images, achieving a better retrieval performance than the original BLIP-2 model.
B.4 Attention Map of CLIP and ELIP-C